OpenStudy (anonymous):

A plane flew 1740 km in 6 hours with a tail wind. On the return trip, going into the wind, the plane flew 570 km in 3 hours. What is the plane's speed if there was no wind? What is the speed of the wind? Write the system of equations that would be used to solve this problem.

OpenStudy (anonymous):

(r+w)6=1740 (r-w)3=570

OpenStudy (anonymous):

(r+w)3=1740 (r-w)6=570

OpenStudy (anonymous):

(r+w)3=1740 (r-w)6=570

OpenStudy (anonymous):

which one?

OpenStudy (anonymous):

The first one

OpenStudy (anonymous):

At the beginning the speed of both the plane (r km/hr) and the wind (km/hr) add up to (r+w) km/hr multiply by the total time spent in hr you get (r+w)3=total distance traveled=1746

OpenStudy (anonymous):

can u help me with a different problem please

OpenStudy (anonymous):

For the second equation, since the plane traveled a distance 570 km, then the speed of the plane must be greater than the speed of the wind and both are in opposite directions then (r-w)km/hr x 3 hr= 570